The sum of the ages of the three Romano brothers is 63. If their ages can be represented as consecutive integers, what is the age of the middle brother?

Answer by ankor@dixie-net.com(22740)

The sum of the ages of the three Romano brothers is 63.
If their ages can be represented as consecutive integers, what is the age of the middle brother?
:
Let x = age of the middle brother
then
(x-1) = age of the younger
and
(x+1) = age of the older
:
x + (x-1) + (x+1) = 63
3x = 63
x = 21 yrs is the middle brother